# Sat2Scene: 3D Urban Scene Generation from Satellite Images with Diffusion

### v1

@@ 1 Introduction
The direct generation of scenes from satellite imagery sparks exciting possibilities for seamlessly integrating diverse real-life environments into games, films, and map services. However, this task remains challenging due to the significant view differences between satellite images and street-view images on the ground, coupled with the expansive scale of the scenes under consideration. Prior efforts primarily focused on image or video generation [22, 32, 29, 18], with relatively limited exploration into scene generation. The salient advantage of scenes resides in their adaptability for rendering images from arbitrary views, which is a feature particularly beneficial for the aforementioned practical applications.
Some works [22, 32, 29] tend to generate street-view images from the geometry estimated based on the satellite height map. Their limitations arise from the individual generation through 2D GANs, leading to the resulting images holding a noticeable lack of inter-frame consistency. Sat2Vid [18] further advances the generation of street-view images with multi-view consistency by explicitly building the 2D-3D correspondence map between pixels in frames and point clouds. Nevertheless, the 3D point cloud is constructed by back-projecting pixels in all 2D frames, restricting its ability to synthesize video along predefined trajectories rather than arbitrary views. Furthermore, memory usage significantly increases with growing trajectory length, imposing constraints on the permissible video length.
Satellite w/ trajectory
Generated ground-view
Generated bird-view
@@
  • We present a novel diffusion-based framework, Sat2Scene, which is able to generate 3D urban scenes just from satellite images.
  • To ensure consistent image generation from any view, a novel diffusion model with sparse representations is proposed to generate scene features tightly associated with the geometry directly in 3D space.a To the best of our knowledge, we are the first to combine diffusion models with 3D sparse representations.
@@ 2 Related work
3D-aware generative models that do not rely on diffusion models have been extensively researched with various scene representations. PixelNeRF [37] extends NeRF [24] in a feed-forward manner without test-time optimization. The sampling points in the rendering ray aggregate features from given views, which are used for color and density estimation. However, due to the less accurate point correspondence between views, the rendered results often exhibit blurriness. Alternatively, some methods integrate GAN settings with neural rendering techniques. They commonly employ intermediate explicit representations associated with simplified space structures or sparse grids to enhance memory efficiency and scalability for scenes of varying sizes. For example, GSN [8] generates a 2D floor plan as the intermediate explicit representation for unconditional indoor living scene generation. Similarly, EG3D [3] adopts a tri-plane representation, contributing to both the efficiency and effectiveness of the generation process. CompNVS [19] tries to solve the 3D inpainting or completion task via conditional generation based on NSVF [21] representation using a sparse grid. GANcraft [12] is designed for large-scale scene generation given sparse geometry with semantic information. The approach leverages a pre-trained SPADE model [28] to generate pseudo ground-truth images, which are then used to perform a NeRF-like optimization with a style controller. Notably, GANcraft’s performance is heavily dependent on the image quality of the 2D generative models and necessitates per-scene training, which is non-feed-forward.
@@
Satellite-to-ground synthesis is a very typical application of cross-view generation, which focuses on synthesizing images from a distinctly different perspective of a provided top-view satellite image. S2G [22] is the first work in this area, considering geometry consistency in satellite-to-ground synthesis. The approach predicts a density voxel grid from the satellite height map and transforms it to ground depth and semantics panorama, followed by a 2D generative model to obtain the street view. Its follow-up Sat2Vid [18] further addresses the temporal inconsistency challenge when generalizing S2G to video synthesis, via maintaining a correspondence map between 3D points and 2D frame pixels. The generator performs directly in 3D space, facilitating the natural maintenance of multi-view consistency in the rendered frames. Nevertheless, the point cloud is pre-defined following the rendering poses hence the size of the point cloud would explode with the length of video sequence increasing. Recently, InfiniCity [20] implements the satellite-to-ground generation on a city-wide scale. However, since it adopts GANcraft [12] as the generative model, the method still requires per-scene training and test-time optimization. Sat2Density [29] places emphasis on top-view density prediction without depth supervision. It achieves this by leveraging the view relationship between satellite and ground, and integrates neural rendering techniques to enhance synthesis quality. But due to the 2D synthesis approach, the resulting videos lack good temporal consistency.
@@ 3.1 Generation
where $t$ is uniformly sampled from $\{1,\ldots,T\}$ with $T$ the number of denoising steps, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\mathbf{C}_{t}$ denotes the noisy version of the color map $\mathbf{C}$ at time step $t$, $\mathbf{m}\in[0,1]^{N}$ is a mask indicating the point confidence with $\otimes$ the broadcasted Hadamard multiplication. The background generation follows the SOTA diffusion model [30], denoting $\boldsymbol{\mathcal{D}}_{B}$, resulting in a panoramic sky image $\mathbf{B}\in[0,1]^{H_{s}\times W_{s}\times 3}$.
@@ 3.3 Implementation details
Training. The entire pipeline is implemented and trained using PyTorch and Minkowski Engine [6]. The foreground denoising network $\boldsymbol{\mathcal{D}}$ is a time-conditional UNet [31], while the feature extractor $\boldsymbol{\mathcal{E}}$ is a vanilla UNet. Both networks adopt the provided default network architectures with 3D sparse voxel convolutions. In the 3D diffusion model, we use $T$$=$1000, and for the rendering phase, the hyper-parameters are set as follows: $D$$=$32, $K$$=$12, $R$$=$3m, and $\beta$$=$0.5. The framework runs on a single NVIDIA Tesla A100 GPU with a memory of 40GB. The generation and rendering phases are trained separately. The training of the 3D diffusion model takes approximately 5 days, while the training of the neural renderer takes around 2 days, both starting from scratch. The network voxelization size is set to 6.67cm and 5cm in the two phases respectively in order to fit the GPU memory. All the code will be made publicly available upon acceptance of the paper.
@@ 4.1 Configuration
Metrics. We evaluate our model and the baselines using the combined validation and test splits of the HoliCity dataset, containing 624 scenes and around 5k single images in total. For quantitative evaluation, we adopt commonly used FID [14] and KID [1] to measure the generated image quality by comparing the distribution similarity between sets of synthesized single frames and real images. We evaluate the temporal consistency across video frames via lower-level metrics PSNR, SSIM, and LPIPS [41], which are only computed in the overlapping areas between neighboring frames. Also, considering that none of the above metrics can simultaneously measure generation quality and inter-frame consistency, we further adopt FVD and KVD [35], which are the video version of FID & KID, being able to focus more on the overall quality. They use video-classification network I3D [2] as the backbone, where the spatiotemporal features are utilized and the temporal relationship can be well modeled. To overcome the insufficient number of test scenes to compute the metrics, during calculation, we pass augmented videos with each frame as the starting frame circularly, as the 8 poses of each scene in the dataset can form a loop. This enhancement ensures that the resulting number of videos is still enough to compute those two metrics. Additionally, we incorporated a user study of 32 people to evaluate the generated videos’ quality, consisting of 30 multiple-choice questions. For each question, the participant was asked to select a single generated video from three options (baselines and ours) that they considered more visually plausible.
@@
- MVDiffusion [34] is a concurrent advancement in multi-view generative models, conditioned on depths and camera poses. The architecture is based on LDMs [30] and ControlNet [40], together with the proposed correspondence-aware attention (CAA) module designed to enhance multi-view consistency. We train the model on the HoliCity dataset by following the proposed two-stage process, i.e., training the single-view depth-conditioned diffusion model backbone first and fixing it, followed by training the ControlNet part with the CAA module for multiple views. We use their default hyper-parameters during training but use a uniform text prompt, "a street with buildings".
@@
- InfiniCity [20] is a recent advancement for the city-scale generation. Since their code is not publicly available but the model is trained with the HoliCity dataset, we directly take the FID and KID scores from their paper for reference only.
@@ 4.3 Qualitative comparison
Fig. 4 includes the visualization results for six samples generated by our method and two baselines. For MVDiffusion [34] and our method, we generate 48 frames for every scene with the interpolated poses between the 8 poses given by the dataset. Since the memory needed by Sat2Vid [18] inflates extremely with a lot of frame generation, we keep its frame number to 8 and decrease its video FPS for synchronized visualization with other 48-frame videos. Both Sat2Vid and our method show good consistency between frames since the texture features are tiled on the scene geometry. Instead, although MVDiffusion links frame texture features via the geometry correspondence and restricts their similarity, it can hardly ensure consistency between frames completely. When considering photorealism, our method and MVDiffusion generate more plausible textures for the building facades than Sat2Vid. MVDiffusion can output more diversified building textures than ours. We infer this is because the LDMs [30] embedded in MVDiffusion have a higher ability to memorize and generate texture patterns with its large number of pre-trained parameters, compared with the diffusion model used by our method. In short, ours achieves a commendable equilibrium between the photorealism and consistency of synthesized results, showcasing a balance superior to Sat2Vid and MVDiffusion.
4.4 Ablation study
w/o pt-rsmp
w/o pt-aggr
w/o dep-sup
Ours
We further conduct ablative experiments aiming at validating the effectiveness of the three components utilized in our method.
@@ 4.4 Ablation study
Satellite
MVDiffusion [34]              Ours
@@
*Refer to captionmedia/omnicity/sat_0.mp4 Refer to caption
[…]
*Refer to captionmedia/omnicity/omnicity_0.mp4
@@
*Refer to captionmedia/omnicity/sat_3.mp4 Refer to caption
[…]
*Refer to captionmedia/omnicity/omnicity_3.mp4
@@ 5 Conclusion
In this supplementary material, we provide additional experiment results (Sec. 6) as well as further details (Sec. 7), and discuss the limitations of our method and potential future directions (Sec. 8).
@@ 6 Additional experiment results
*Refer to captionmedia/supp_baseline/supp_MVDiff_6res.mp4
@@
*Refer to captionmedia/supp_baseline/supp_Ours_6res.mp4
@@
w/o pt-rsmp
w/o pt-aggr
w/o dep-sup
Ours
Visualization of the denoising process can be seen in Fig. 7. The first few seconds of the videos clearly demonstrate the process of the texture changing from noise to meaningful visual patterns. We observe larger noise patches in the background and its denoising process holds good stability since it uses LDMs [30].
@@
Additional qualitative baseline comparisons are provided in Fig. 8 as the extension of Fig. 4.
An additional qualitative ablation study is provided in Fig. 9 as the extension of Fig. 5.
7 Additional details
OmniCity [17] video poses. All the videos in Fig. 1, Fig. 6, and Fig. 7 use trajectories along the right side of the road. The ground-view poses are at a height level of 2m, ${}^{\circ}$ pitch facing the sky and zero roll, while bird-view poses are at a height level of 10m, ${}^{\circ}$ pitch facing the road and zero roll.
User study. We also provide the offline version of the questionnaire form used for our user study, please see the attachment for reference.
@@ 8 Limitations and discussion
  • Our model may not handle very large-scale scenes, e.g., a city-scale scene due to potentially limited computation resources. Generating blocks separately can be an alternative to solve the scale problem, but it also introduces the potential problem of texture discontinuity between neighboring blocks.
@@
  • Our model only performs on the potentially visible surfaces rather than all the surfaces of the scene to reduce computational efforts. The invisible grounds at the base of the buildings were also removed.
  • The generated building textures do not hold very well diversity, which may be because the size of the 3D networks cannot be as large as the 2D networks, which limits its capacity for diverse generations. On the other hand, the details level of the ground truth geometry may not be high enough. For example, it does not reflect the slightly higher sidewalks or recessed windows on the building facades, which could also make learning difficult.
  • The road surface is not very well generated. It may also be due to the dataset issue, where the roads are often occluded by cars or passengers. Also, the GT poses are often looking a bit up to the sky so the road surfaces may be less covered and trained.
@@
Below we discuss potential future development directions. Instead of the current setting which is generating the whole scene, we can further divide the whole geometry into small elements, e.g., building instances, and generate textures individually, which could lead to better diversity. In terms of network architecture, we may also follow LDMs [30] to perform the diffusion models in a latent space, given that there is enough 3D urban scene data to train a good auto-encoder with a sparse setting. Furthermore, we can incorporate satellite image information mainly for road surface generation. Also, our model can be combined with a natural language setting, or can even build a side branch similar to ControlNet [40] on top of the trained one, to further extend the model to a conditional generation, e.g., having semantic information as input.

### v2

@@ 1 Introduction
The direct generation of scenes from satellite imagery sparks exciting possibilities for seamlessly integrating diverse real-life environments into games, films, and map services. However, this task remains challenging due to the significant view differences between satellite images and street-view images on the ground, coupled with the expansive scale of the scenes under consideration. Prior efforts primarily focused on image or video generation [23, 34, 31, 19], with relatively limited exploration into scene generation. The salient advantage of scenes resides in their adaptability for rendering images from arbitrary views, which is a feature particularly beneficial for the aforementioned practical applications. Conventional cross-view synthesis works [23, 34, 31] tend to generate street-view images from the geometry estimated based on the satellite height map. Their limitations arise from the individual generation through 2D GANs, leading to the resulting images holding a noticeable lack of inter-frame consistency. Sat2Vid [19] further advances the generation of street-view images with multi-view consistency by explicitly building the 2D-3D correspondence map between pixels in frames and point clouds. Nevertheless, the 3D point cloud is constructed by back-projecting pixels in all 2D frames, restricting its ability to synthesize video along predefined trajectories rather than arbitrary views. Furthermore, memory usage significantly increases with growing trajectory length, imposing constraints on the permissible video length.
Satellite w/ trajectory Generated ground-view Generated bird-view
[…]
@@
  • We present a novel diffusion-based framework Sat2Scene for direct 3D scene generation, which is able to generate 3D urban scenes just from satellite images.
  • To ensure consistent image generation from any view, a novel diffusion model with sparse representations is proposed to generate scene features tightly associated with the geometry directly in 3D space. To the best of our knowledge, we are the first to combine diffusion models with 3D sparse representations.
@@ 2 Related work
3D-aware generative models that do not rely on diffusion models have been extensively researched with various scene representations. PixelNeRF [40] extends NeRF [26] in a feed-forward manner without test-time optimization. The sampling points in the rendering ray aggregate features from given views, which are used for color and density estimation. However, due to the less accurate point correspondence between views, the rendered results often exhibit blurriness. Alternatively, some methods integrate GAN settings with neural rendering techniques. They commonly employ intermediate explicit representations associated with simplified space structures or sparse grids to enhance memory efficiency and scalability for scenes of varying sizes. For example, GSN [9] generates a 2D floor plan as the intermediate explicit representation for unconditional indoor living scene generation. Similarly, EG3D [3] adopts a tri-plane representation, contributing to both the efficiency and effectiveness of the generation process. CompNVS [20] tries to solve the 3D inpainting or completion task via conditional generation based on NSVF [22] representation using a sparse grid. GANcraft [13] is designed for large-scale scene generation given sparse geometry with semantic information. The approach leverages a pre-trained SPADE model [30] to generate pseudo ground-truth images, which are then used to perform a NeRF-like optimization with a style controller. Notably, GANcraft’s performance is heavily dependent on the image quality of the 2D generator and necessitates per-scene training, which is non-feed-forward.
@@
Satellite-to-ground synthesis is a very typical application of cross-view generation, which focuses on synthesizing images from a distinctly different perspective of a provided top-view satellite image. S2G [23] is the first work in this area, considering geometry consistency in satellite-to-ground synthesis. The approach predicts a density voxel grid from the satellite height map and transforms it to ground depth and semantics panorama, followed by a 2D generative model to obtain the street view. Its follow-up Sat2Vid [19] further addresses the temporal inconsistency challenge when generalizing S2G to video synthesis, via maintaining a correspondence map between 3D points and 2D frame pixels. The generator performs directly in 3D space, facilitating the natural maintenance of multi-view consistency in the rendered frames, showing better plausibility than 2D video-to-video translation advances Vid2Vid [38] and WC-Vid2Vid [25]. Nevertheless, the point cloud is pre-defined following the rendering poses hence the size of the point cloud would explode with the length of the video sequence increasing. Recently, InfiniCity [21] implements the satellite-to-ground generation on a city-wide scale. However, since it adopts GANcraft [13] as the generative model, the method still requires per-scene training and test-time optimization. Sat2Density [31] places emphasis on top-view density prediction without depth supervision. It achieves this by leveraging the view relationship between satellite and ground, and integrating neural rendering techniques to enhance synthesis quality. But due to this, the resulting videos lack good temporal consistency.
@@ 3.1 Generation
where $t$ is uniformly sampled from $\{1,\ldots,T\}$ with $T$ the number of denoising steps, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\mathbf{C}_{t}$ denotes the noisy version of the color map $\mathbf{C}$ at time step $t$, $\mathbf{m}\in[0,1]^{N}$ is a mask indicating the point confidence with $\otimes$ the broadcasted Hadamard multiplication. The background generation follows the SOTA diffusion model [32], denoting $\boldsymbol{\mathcal{D}}_{B}$, resulting in a panoramic sky image $\mathbf{B}\in[0,1]^{H_{B}\times W_{B}\times 3}$.
@@ 3.3 Implementation details
Training. The entire pipeline is implemented and trained using PyTorch and Minkowski Engine [6]. The foreground denoising network $\boldsymbol{\mathcal{D}}$ is a time-conditional UNet [33], while the feature extractor $\boldsymbol{\mathcal{E}}$ is a vanilla UNet. Both networks adopt the provided default network architectures with 3D sparse voxel convolutions. In the 3D diffusion model, we use $T$$=$1000, and for the rendering phase, the hyper-parameters are set as follows: $D$$=$32, $K$$=$12, $R$$=$3m, and $\beta$$=$0.5. The framework runs on a single NVIDIA Tesla A100 GPU with a memory of 40GB. The generation and rendering phases are trained separately. The training of the 3D diffusion model takes approximately 5 days, while the training of the neural renderer takes around 2 days, both starting from scratch. The network voxelization size is set to 6.67cm and 5cm in the two phases respectively in order to fit the GPU memory.
@@ 4.1 Configuration
Metrics. We evaluate our model and the baselines using the combined validation and test splits of the HoliCity dataset, containing 624 scenes and around 5k single images in total. For quantitative evaluation, we adopt commonly used FID [15] and KID [1] to measure the generated image quality by comparing the distribution similarity between sets of synthesized single frames and real images. We evaluate the temporal consistency across video frames via lower-level metrics PSNR, SSIM, and LPIPS [44], which are only computed in the overlapping areas between neighboring frames. Also, considering that none of the above metrics can simultaneously measure generation quality and inter-frame consistency, we further adopt FVD and KVD [37], which are the video version of FID & KID, being able to focus more on the overall quality. They use video-classification network I3D [2] as the backbone, where the spatiotemporal features are utilized and the temporal relationship can be well modeled. To overcome the insufficient number of test scenes to compute the metrics, during calculation, we pass augmented videos with each frame as the starting frame circularly, as the 8 poses of each scene in the dataset can form a loop. This enhancement ensures that the resulting number of videos is still enough to compute those two metrics. Additionally, we incorporated a user study of 32 people to evaluate the quality of generated videos, consisting of 30 multiple-choice questions. For each question, the participant was asked to select a single generated video from three options (baselines and ours) that they considered more visually plausible.
@@
- MVDiffusion [36] is a concurrent advancement in multi-view generative models, conditioned on depths and camera poses. The architecture is based on LDMs [32] and ControlNet [43], together with the proposed correspondence-aware attention (CAA) module designed to enhance multi-view consistency. We train the model on the HoliCity dataset by following the proposed two-stage process, i.e., training the single-view depth-conditioned diffusion model backbone first and fixing it, followed by training the ControlNet part with the CAA module for multiple views. We use their default hyper-parameters during training but use a uniform text prompt, “a street with buildings”.
@@
- InfiniCity [21] is a recent advancement for the city-scale generation based on GANcraft [13]. Since their code is not publicly available but the model is trained with the HoliCity [45] dataset, we directly take the FID and KID scores from their paper for reference only.
@@ 4.3 Qualitative comparison
Fig. 4 includes the visualization results for six samples generated by our method and two baselines. For MVDiffusion [36] and our method, we generate 48 frames for every scene with the interpolated poses between the 8 poses given by the dataset. Since the memory needed by Sat2Vid [19] inflates extremely with a lot of frame generation, we keep its frame number to 8 and decrease its video FPS for synchronized visualization with other 48-frame videos. Both Sat2Vid and our method show good consistency between frames since the texture features are tiled on the scene geometry. Instead, while MVDiffusion connects frame texture features via geometric correspondences and limits the similarity, it may still struggle to guarantee complete consistency between frames. This challenge could become particularly severe for 2D methods when dealing with a smaller overlap ratio in the training data. For instance, the average foreground overlap between two neighboring “key” frames is $\sim$42in HoliCity [45], as opposed to $\sim$65in ScanNet [8] used in MVDiffusion, which could be the primary reason that it presents an inferior consistency compared with the results reported in its original paper. When considering photorealism, our method and MVDiffusion generate more plausible textures for the building facades than Sat2Vid. MVDiffusion can output more diversified building textures than ours. We infer this is because the LDMs [32] embedded in MVDiffusion have a higher ability to memorize and generate texture patterns with its large number of pre-trained parameters, compared with the diffusion model used by our method. In short, ours achieves a commendable equilibrium between the photorealism and consistency of synthesized results, showcasing a balance superior to Sat2Vid and MVDiffusion.
@@
4.4 Ablation study
w/o pt-rsmp w/o pt-aggr w/o dep-sup Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
We further conduct ablative experiments to validate the effectiveness of the three components utilized in our method.
@@ 4.4 Ablation study
Satellite MVDiffusion [36]              Ours
@@ 5 Conclusion
In this supplementary material, we provide further details (Sec. 6), present additional experiment results (Sec. 7), and discuss the limitations of our method along with potential future directions (Sec. 8).
6 Additional details
OmniCity [18] video poses. All the videos in Fig. 1, Fig. 6, and Fig. 7 use trajectories along the right side of the road. The ground-view poses are at a height level of 2m, 15 pitch facing the sky and zero roll, while bird-view poses are at a height level of 10m, 15 pitch facing the road and zero roll.
Other related baselines. We did not include experiments with Vid2Vid [38], WC-Vid2Vid [25] and Sat2Density [31] due to the following reasons. Vid2Vid [38] has already exhibited deficiencies in temporal consistency, as observed in Sat2Vid [19]. WC-Vid2Vid [25], on the other hand, necessitates semantic video input, which exceeds the scope of our problem setting. Furthermore, Sat2Density [31] requires for precise satellite-ground image correspondence, which is unavailable in the HoliCity [45] dataset.
@@ 7 Additional experiment results
Visualization of the denoising processes on the exemplary scenes are presented in Fig. 7. In the initial seconds of the videos, the transformation of texture from complete noise to meaningful visual patterns is evident. Larger noise patches are observed in the sky background, and the denoising process exhibits stability due to the utilization of LDMs [32].
@@
Supplementary experiments including additional qualitative baseline comparisons and a qualitative ablation study, are presented in Fig. 8 and Fig. 9, respectively. These figures serve as extensions of Fig. 4 and Fig. 5.
The inferior consistency of MVDiffusion [36] compared to its original paper is evident. In addition to the overlap ratio mentioned in Sec. 4.3, we believe this may also be attributed to factors such as text prompt and depth detail. MVDiffusion [36] utilizes long and diversified text prompts to provide the network with rich scene information. To ensure a fair comparison with our approach and other baselines, we use a universal text prompt, which may increase the difficulty of learning color information or detailed texture. As for the depth detail, the indoor depths in MVDiffusion [36] exhibit high quality, even allowing for the recognition and inference of object boundaries. In contrast, the outdoor depths in HoliCity [45] appear more “flat” and lack high-frequency details such as windows and facade decorations. Thus, learning the mapping from geometry to appearance in the HoliCity dataset becomes more challenging compared to the indoor scenarios in MVDiffusion [36].
@@ 8 Limitations and discussion
Large-scale generation. Our model may not handle very large-scale scenes, e.g., a city-scale scene due to potentially limited computation resources. Generating blocks separately can be an alternative to solve the scale problem, but it also introduces the potential problem of texture discontinuity between neighboring blocks.
@@
Computation only on surfaces. Our model only performs on the potentially visible surfaces rather than all the surfaces of the scene to reduce computational efforts. The invisible grounds at the base of the buildings were also removed.
Style diversity. The generated building textures do not hold very well diversity, which can be attributed to several factors: (1) The GT geometry in HoliCity [45] dataset is not sufficiently detailed, lacking representations of subtle features such as slightly elevated sidewalks or recessed windows on building facades. Such a limitation can hinder the learning process for mapping geometry to appearance. (2) Our approach generates entire scenes simultaneously, necessitating the learning of appearance correlations among different building instances, in addition to individual building’s appearance. This may lead the network to pay less attention to diversity. (3) 3D networks typically have a smaller network scale (number of layers, channels, etc.) than 2D ones due to 3D convolution, which limits their capacity to generate diverse appearances. As potential solutions, one can consider introducing latent diffusion models in 3D sparse space or adopting generation based on individual building instances. Also, incorporating semantics as input conditions may contribute to enhancing diversity.
Road surface. The road surface is not very well generated and there are three potential reasons stemming from the dataset: (1) The presence of cars, passengers, and shadows on roads in GT images poses a challenge to the segmentation model in filtering them out, potentially resulting in black spots on the road surface during inference. (2) Due to the GT poses often looking a bit up to the sky, a small area around the camera origin tends to lack sufficient supervision signals during training. (3) The pixel ratio for buildings ($\sim$43.2%) surpasses that of roads ($\sim$15.8%) in GT images, leading the network to prioritize learning building facades in this unbalanced scenario. A potential solution for addressing (3) is to introduce a balance factor in the loss functions or employ distinct networks dedicated to handling buildings and roads, respectively.
@@
Future direction. Below we discuss potential future development directions. Instead of the current setting which is generating the whole scene, we can further divide the whole geometry into small elements, e.g., building instances, and generate textures individually, which could lead to better diversity. In terms of network architecture, we may also follow LDMs [32] to perform the diffusion models in a latent space, given that there is enough 3D urban scene data to train a good auto-encoder with a sparse setting. Furthermore, we can incorporate satellite image information mainly for road surface generation. Also, our model can be combined with a natural language setting, or can even build a side branch similar to ControlNet [43] on top of the trained one, to further extend the model to a conditional generation, e.g., having semantic information as input.
w/o pt-rsmp w/o pt-aggr w/o dep-sup Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption